OpenStudy (anonymous):

solve this equation ln(2x+5)+ln(x-4)-2lnx=0

OpenStudy (aum):

log(A) + log(B) - log(C) = log(AB/C). Write 0 as log(1) Equate the two and solve for x. Note if log(P) = log(Q) then P = Q.

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

what does x equal rounded to 2 decimal places

OpenStudy (aum):

\[ \ln(2x+5) + \ln(x-4) - 2\ln(x) = 0 \\ \ln\{(2x+5) * (x-4)\} - \ln(x^2) = 0 \\ \ln\{(2x+5) * (x-4)/x^2\} = 0 \\ \ln\{(2x+5) * (x-4)/x^2\} = \ln(1) \\ \{(2x+5) * (x-4)/x^2\} = 1 \\ (2x+5)(x-4) = x^2 \]Solve for x

OpenStudy (anonymous):

i dont think this is right

OpenStudy (anonymous):

i need to end up with 1 answer

OpenStudy (anonymous):

look right to me

OpenStudy (aum):

When you solve the quadratic you will get 2 values for x. But you have to test each value in the original problem to make sure we don't end up taking log of a number less than or equal to 0. If so discard that x value as an extraneous solution.

OpenStudy (anonymous):

i got root 6 and i root 6 over 2

OpenStudy (anonymous):

im supposed to have a decimal

OpenStudy (aum):

(2x+5)(x-4) = x^2 2x^2 - 8x + 5x - 20 = x^2 x^2 - 3x - 20 = 0 x = [ 3 +/- sqrt{9 + 80} ] / 2 x = 1/2 * (3 + sqrt(89)} or x = 1/2 * (3 - sqrt(89)} The second one is negative and we can discard it because we can't take log of a negative number. So x = 1/2 * (3 + sqrt(89)} = 6.22 (to two decimal places).